OpenStudy (anonymous):

2(h – 8) – h = h – 16 solve

OpenStudy (anonymous):

2h - 16 -h = h -16 h - 16 = h -16 2h = -16 + 16 2h = 0 h=0

OpenStudy (anonymous):

h can be all real number

OpenStudy (anonymous):

\[h \in \mathbb{R}\]